# Charm associated production of the pseudoscalar Higgs boson $h^{\prime\prime}$ in the MCPM’ at the LHC

### v1

@@ Abstract
We discuss charm associated production of the pseudoscalar Higgs boson $h^{\prime\prime}$ in the MCPM’. In our analysis we assume $m_{h^{\prime\prime}}$ = 95.4 GeV, which corresponds to an enhancement observed by the CMS collaboration in the $\gamma\gamma$ channel. As discussed recently, the MCPM’ is consistent with the CMS enhancement. In this model the $h^{\prime\prime}$ Higgs boson of this mass decays dominantly into $c\bar{c}$ jets. The cross section for the associated $h^{\prime\prime}$ production in $pp\to h^{\prime\prime}c\bar{c}$ is found to be almost as big as that for the Drell-Yan (DY) $c\bar{c}\to h^{\prime\prime}$ production, much larger than that for two-gluon fusion $gg\to h^{\prime\prime}$. Since in the MCPM’ the dominant $h^{\prime\prime}$ decay is into the $c\bar{c}$ channel we have to deal with the $pp\to c\bar{c}c\bar{c}$ process with $c$ and $\bar{c}$ jets. In addition to the signal processes, we have sizeable single parton scattering (SPS) and double parton scattering (DPS) nonresonant background. In our analysis we include diagrams of gluon-gluon fusion $gg\to h^{\prime\prime}$ which lead to the $c\bar{c}c\bar{c}$ final state. We discuss also the possible role of the reaction $pp\to h^{\prime\prime}h^{\prime\prime}$. Assuming 100% c and $\bar{c}$ jet tagging we discuss in detail how to improve the signal-to-background ratio. Our analysis shows promising resonance-like enhancements and relatively large cross sections in the $c\bar{c}c\bar{c}$ channel within the MCPM’. The ALICE-3 project would be ideal in the future to perform the experimental studies corresponding to the theoretical studies discussed here.
@@ 1 Introduction
In this paper we are studying consequences of the model MCPM’, which was introduced in Maniatis:2023aww. The MCPM’ is a two-Higgs-doublet model (2HDM) that, in our view, exhibits several noteworthy features when compared with experimental observations. The basis of the MCPM’ is the maximally CP symmetric two-Higgs-doublet model, MCPM, which was introduced in Maniatis:2007de and a short description of it is given in Maniatis:2010sb. The MCPM has four generalized CP (GCP) symmetries: see Table 1 of Maniatis:2023aww. These symmetries lead to highly constrained Yukawa interactions. In the strict symmetry limit only the third generation fermions, $\tau$, $t$, $b$, are massive and they couple to the first Higgs doublet $\varphi_{1}$. The fermions $\mu$, $c$, $s$ of the second family are massless and couple to the second Higgs doublet $\varphi_{2}$. The fermions $e$, $u$, $d$ are uncoupled to the Higgs bosons. Phenomenological predictions/estimates in the MCPM have been worked out in Maniatis:2007de; Maniatis:2009vp; Maniatis:2009by; Maniatis:2010sb; Brehmer:2012hh. However, clearly the highly symmetric MCPM can only be an approximate representation of what is observed in nature. Therefore, in Maniatis:2023aww a realistic model, the MCPM’, was constructed. The key idea there was to generate masses for fermions of the second and first family by effective interactions originating from integrating out high-mass particles having suitable couplings. In this way, masses for all fermions except neutrinos were generated. That is, in the MCPM’ we work with massless neutrinos which should be alright for the reactions which we want to study. In addition, electroweak symmetry breaking (EWSB) and a realistic, nontrivial, Cabibbo-Kobayashi-Maskawa (CKM) matrix are generated in the MCPM’.
@@
The $\rho^{\prime}$ behaves very much like the Standard-Model (SM) Higgs boson. Therefore, we set ParticleDataGroup:2024cfk
@@
The Higgs bosons $h^{\prime}$ and $h^{\prime\prime}$ have a scalar and pseudoscalar coupling to fermions, respectively. The main coupling of $h^{\prime}$ and $h^{\prime\prime}$ is to charm quarks; see (A.81) of Maniatis:2023aww. The couplings of the Higgs bosons of the MCPM’ among themselves and to gauge bosons are as in the MCPM and are listed in appendix A of Maniatis:2009vp. The Yukawa couplings of the Higgs bosons in the MCPM’ are given in appendix A of Maniatis:2023aww; see (A50)–(A58) and (A81) there.
In Maniatis:2023aww predictions of the MCPM’ were worked out concerning the following reaction at the LHC:
@@
The comparison with the CMS results CMS:2023yay, where a possible $\gamma\gamma$ signal at 95.4 GeV was reported, is quite encouraging. Therefore, in Maniatis:2023aww tentatively the mass of the $h^{\prime\prime}$ boson was set to
@@
In Maniatis:2023aww also the leptonic decays of the mesons $D^{\pm}$ and $D_{s}^{\pm}$ were studied. It was shown that the contributions of the charged Higgs bosons $H^{\pm}$ of the MCPM’ lead to deviations from the predictions of the Standard Model (SM). From comparison with experiment it was concluded that the $H^{\pm}$ mass $m_{H^{\pm}}$ should be around 300 GeV.
@@
In the reaction (6) we have two charm and two anticharm jets. Due to the resonance $h^{\prime\prime}$ we expect that in experiment one should find an enhancement in the $c\bar{c}$ mass distribution at $m_{h^{\prime\prime}}$. Such a study would profit enormously if charm tagging can be realized, a challenging task at the LHC. However, it has been shown to be feasible. ATLAS, for instance, reports a charm-tagging efficiency of about 30% ATLAS:2022qxm.
@@ 2 Results: generalities and integrated cross sections
We want to show our results for the calculation of charm-quark production, $c\bar{c}c\bar{c}$, in proton–proton collisions at the LHC. We compare the predictions of the MCPM’ versus those of the SM. We employ the MadGraph5 Monte-Carlo event generator to simulate the events Alwall:2011uj. The MCPM has been implemented in MadGraph5 Brehmer:2012hh. We extend this implementation by adding the effective gluon-gluon-$h^{\prime\prime}$ interaction. The $h^{\prime\prime}$ boson coupling to the gluons (and photons) is only possible via loops. The calculation of this effective interaction is presented in Appendix A.
For the process of associated $h^{\prime\prime}$ charm production we adopt for the center-of-mass energy the current LHC value of $\sqrt{s}=$ 13.6 TeV. The mass of the pseudoscalar $h^{\prime\prime}$ boson is chosen to match the possible resonance observed at CMS, that is, $m_{h^{\prime\prime}}=95.4$ GeV, (5). The total width of the $h^{\prime\prime}$ boson has been calculated in Maniatis:2009vp giving, for the chosen mass value
@@
For the parton densities we use the set MSHT20lo_\!\_\!as130 Bailey:2020ooq, the factorization and the renormalization scales are set to $\mu_{F}=\mu_{R}=95.4$ GeV.
@@ 3 Distributions
Now we come to the distributions for the reactions (6) and (7). The $c$ and $\bar{c}$ quarks in these reactions will manifest themselves as $c$ and $\bar{c}$ jets. Therefore, we will now specify the cuts which we apply at parton level in order to obtain distributions which should be also valid at the jet level. In the following we assume that the $c$ and $\bar{c}$ jets are identified by flavor-tagging. We are aware of the fact that this is a tough requirement for experimental analyses. But, as already mentioned, prospects to achieve this are quite good ATLAS:2022qxm; CMS-DP-2023-006.
@@ 3.2 Cross sections for $p+p\to h^{\prime\prime}+c+\bar{c}+X$ with $h^{\prime\prime}\to c+\bar{c}$
The $h^{\prime\prime}$ boson couples in particular to the charm quarks with a coupling strength proportional to the top-quark mass. This is a special feature of the MCPM’ - eventually originating from the underlying CP symmetry; see Maniatis:2023aww. With the application of the cuts as outlined in Section 3.1 we find a total cross section of
@@ 3.3 Cross sections for $p+p\to c+\bar{c}+c+\bar{c}+X$
Figure 9 presents a comparison of SM calculations for $d\sigma/dM_{c\bar{c}}$ obtained using different theoretical approaches. The production of the $c\bar{c}c\bar{c}$ final state within the Standard Model has been thoroughly analyzed in Maciula:2013kd; vanHameren:2014ava; vanHameren:2015wva, where both single-parton scattering (SPS) and double-parton scattering (DPS) mechanisms were considered. In the present study, we adopt the same theoretical frameworks and computational setup. The corresponding numerical results here are obtained using the KaTie Monte Carlo event generator vanHameren:2016kkz, employing both the collinear and $k_{T}$-factorization approaches [^1] We have checked numerically, that the results from MadGraph are consistent with those obtained using the KaTie generator in the collinear mode.. For the unintegrated (transverse momentum dependent) parton distribution functions (uPDFs), we use the PB-LO-2020-set1 model Jung:2025mtd, obtained within the parton-branching method Hautmann:2017fcj, as well as the widely adopted Kimber-Martin-Ryskin (KMR) model Kimber:2001sc; Watt:2003mx, constructed with the strong-ordering condition and by using the MSHT20lo_\!\_\!as130 collinear PDFs as input. These two models can be reliably used in the probed kinematical domain, since they are based on the DGLAP evolution equations. The use of distributions specifically designed for small values of $x$ (as is the case in most of the literature) may not be fully justified in this region. In the case of DPS, we follow a commonly used uncorrelated factorized ansatz Gaunt:2009re; Diehl:2011yj, taking as effective cross section $\sigma_{\mathrm{eff}}=15$ mb. This is a value extracted from experimental analyses of various processes at different kinematic configurations (for recent results see e.g. ATLAS:2016rnd; LHCb:2016wuo; CMS:2013huw; CMS:2017han). The $k_{T}$-factorization approach with the PB-LO-2020-set1 uPDFs yields smaller cross sections in the signal region compared to the collinear calculations (see Fig. 9(a)). In contrast, the calculations performed within the KMR uPDFs yield higher cross sections and result in distributions that are very close to those obtained in the collinear approach (see Fig. 9(b)). Moreover, as evident from the figure, the contribution from the double parton scattering (DPS) mechanism is significantly lower than that of the standard single parton scattering (SPS) process. In the signal region the DPS effect is around 15% of the SPS cross section.
@@ 4 Conclusions
In the present paper we have worked out further consequences of the model MCPM’ for experiments at the LHC. The MCPM’ is a two-Higgs-doublet model (2HDM) which has been introduced in Maniatis:2023aww. The Higgs sector of the MCPM’contains, as in any 2HDM, three neutral Higgs bosons ($\rho^{\prime},h^{\prime},h^{\prime\prime}$) and a charged pair ($H^{\pm}$). The CP-symmetry relations forming the basis of the MCPM’ fix the couplings of these Higgs bosons to the fermions to a large extent. The $\rho^{\prime}$ couples to the fermions like the SM Higgs particle. The bosons $h^{\prime}$ and $h^{\prime\prime}$ couple mainly to charm quarks in a scalar and pseudoscalar way, respectively. The corresponding coupling constants are, however, proportional to the top-quark mass.
In Maniatis:2023aww the MCPM’ was compared to the results for $\gamma\gamma$ production from CMS:2023yay where a possible enhancement at $M_{\gamma\gamma}=95.4$ GeV was reported. Choosing the mass of the pseudoscalar boson $h^{\prime\prime}$ of the MCPM’ at this value compatibility of the MCPM’ with the results of CMS:2023yay was found.
@@
On the experimental side the identification of $c$ and $\bar{c}$ jets is improving (see LHCb:2015tna; Vertesi:2020gqr; CMS:2017wtu). For example, the LHCb collaboration has measured the cross section for $c\bar{c}$ dijets LHCb:2020frr (see also Maciula:2022bfv). The CMS and ATLAS would need to perform their own analyses including jet missidentification probabilities. In our opinion the ALICE-3 project ALICE:2022wwr would be ideal in the future to perform experimental studies corresponding to our theoretical studies discussed here.
@@ Acknowledgements.
We now determine the effective $h^{\prime\prime}$–gluon–gluon interaction. This interaction has no tree-level contribution but arises from a triangle quark loop. Based on this interaction, we get in particular contributions to the process $p+p\to h^{\prime\prime}+c+\bar{c}+X$ via gluon–gluon fusion; see Fig. 2. We estimate that for a mass $m_{h^{\prime\prime}}\sim 95.4\penalty 10000\ \mathrm{GeV}$ in proton–proton collisions, with a center-of-mass energy of $13.6\penalty 10000\ \mathrm{TeV}$, the corresponding Bjorken-$x$ for the gluons is about $0.007$. Gluon densities in the protons are known to be very high at low Bjorken-$x$. On the other hand, the coupling of a gluon pair to the $h^{\prime\prime}$ boson is loop suppressed, with the dominant contribution arising from the charm-quark loop. In the following, we use the $\gamma$-matrix conventions of Bjorken:1965zz and the Levi-Civita symbol with the normalization $\varepsilon_{0123}=1$.
@@
At one loop level we have in the MCPM’ for (A.1) the diagrams shown in Fig. 19. The evaluation of the diagrams of Fig. 19 is straightforward using the $h^{\prime\prime}q\bar{q}$ couplings from (A.81) of Appendix A of Maniatis:2023aww:
@@
With the mass values for the quarks listed in ParticleDataGroup:2024cfk and $\alpha_{s}=0.12$, we find numerically:
@@
Appendix B Digression on total cross sections
Now we want to calculate the total cross section for $h^{\prime\prime}$ production, composed of the cross sections for the Drell–Yan process with $c\bar{c}$ fusion and the $GG$ fusion process discussed in (37) and (38) of Maniatis:2023aww and the process (6) here. For this we shall first make some general considerations.

### v2

@@ Abstract
We discuss charm associated production of the pseudoscalar Higgs boson $h^{\prime\prime}$ in the MCPM’. In our analysis we assume $m_{h^{\prime\prime}}$ = 95.4 GeV, which corresponds to an enhancement observed by the CMS collaboration in the $\gamma\gamma$ channel. As discussed recently, the MCPM’ is consistent with the CMS enhancement. In this model the $h^{\prime\prime}$ Higgs boson of this mass decays dominantly into $c\bar{c}$ jets. The cross section for the associated $h^{\prime\prime}$ production in $pp\to h^{\prime\prime}c\bar{c}$ is found to be almost as big as that for the Drell-Yan (DY) $c\bar{c}\to h^{\prime\prime}$ production, much larger than that for two-gluon fusion $gg\to h^{\prime\prime}$. Since in the MCPM’ the dominant $h^{\prime\prime}$ decay is into the $c\bar{c}$ channel we have to deal with the $pp\to c\bar{c}c\bar{c}$ process with $c$ and $\bar{c}$ jets. In addition to the signal processes, we have sizeable single parton scattering (SPS) and double parton scattering (DPS) nonresonant background. In our analysis we include diagrams of gluon-gluon fusion $gg\to h^{\prime\prime}$ which lead to the $c\bar{c}c\bar{c}$ final state. We discuss also the possible role of the reaction $pp\to h^{\prime\prime}h^{\prime\prime}$. Assuming 100% c and $\bar{c}$ jet tagging we discuss in detail how to improve the signal-to-background ratio. Our analysis shows promising resonance-like enhancements and relatively large cross sections in the $c\bar{c}c\bar{c}$ channel within the MCPM’. In an appendix we discuss the total cross section for $h^{\prime\prime}$ production at the LHC. We also compare distributions for the reactions $pp\to h^{\prime\prime}X$ and $pp\to ZX$ with $h^{\prime\prime}$ and $Z$ decaying to $c\bar{c}$ and $\mu^{+}\mu^{-}$. We hope that the LHC collaborations will in the future perform the experimental studies corresponding to the theoretical studies discussed here.
@@ 1 Introduction
In this paper we are studying consequences of the model MCPM’, which was introduced in Maniatis and Nachtmann (2025). The MCPM’ is a two-Higgs-doublet model (2HDM) that, in our view, exhibits several noteworthy features when compared with experimental observations. The basis of the MCPM’ is the maximally CP symmetric two-Higgs-doublet model, MCPM, which was introduced in Maniatis et al. (2008) and a short description of it is given in Maniatis et al. (2010). The MCPM has four generalized CP (GCP) symmetries: see Table 1 of Maniatis and Nachtmann (2025). These symmetries lead to highly constrained Yukawa interactions. In the strict symmetry limit only the third generation fermions, $\tau$, $t$, $b$, are massive and they couple to the first Higgs doublet $\varphi_{1}$. The fermions $\mu$, $c$, $s$ of the second family are massless and couple to the second Higgs doublet $\varphi_{2}$. The fermions $e$, $u$, $d$ are uncoupled to the Higgs bosons. Phenomenological predictions/estimates in the MCPM have been worked out in Maniatis et al. (2008); Maniatis and Nachtmann (2009, 2010); Maniatis et al. (2010); Brehmer et al. (2013). However, clearly the highly symmetric MCPM can only be an approximate representation of what is observed in nature. Therefore, in Maniatis and Nachtmann (2025) a realistic model, the MCPM’, was constructed. The key idea there was to generate masses for fermions of the second and first family by effective interactions originating from integrating out high-mass particles having suitable couplings. In this way, masses for all fermions except neutrinos were generated. That is, in the MCPM’ we work with massless neutrinos which should be alright for the reactions which we want to study. In addition, electroweak symmetry breaking (EWSB) and a realistic, nontrivial, Cabibbo-Kobayashi-Maskawa (CKM) matrix are generated in the MCPM’.
@@
The $\rho^{\prime}$ behaves very much like the Standard-Model (SM) Higgs boson. Therefore, we set Navas and others (2024)
@@
The Higgs bosons $h^{\prime}$ and $h^{\prime\prime}$ have a scalar and pseudoscalar coupling to fermions, respectively. The main coupling of $h^{\prime}$ and $h^{\prime\prime}$ is to charm quarks; see (A.81) of Maniatis and Nachtmann (2025). The couplings of the Higgs bosons of the MCPM’ among themselves and to gauge bosons are as in the MCPM and are listed in appendix A of Maniatis and Nachtmann (2009). The Yukawa couplings of the Higgs bosons in the MCPM’ are given in appendix A of Maniatis and Nachtmann (2025); see (A50)–(A58) and (A81) there.
In Maniatis and Nachtmann (2025) predictions of the MCPM’ were worked out concerning the following reaction at the LHC:
@@
The comparison with the CMS results 30, where a possible $\gamma\gamma$ signal at 95.4 GeV was reported, is quite encouraging. Therefore, in Maniatis and Nachtmann (2025) tentatively the mass of the $h^{\prime\prime}$ boson was set to
@@
In Maniatis and Nachtmann (2025) also the leptonic decays of the mesons $D^{\pm}$ and $D_{s}^{\pm}$ were studied. It was shown that the contributions of the charged Higgs bosons $H^{\pm}$ of the MCPM’ lead to deviations from the predictions of the Standard Model (SM). From comparison with experiment it was concluded that the $H^{\pm}$ mass $m_{H^{\pm}}$ should be around 300 GeV.
@@
In the reaction (6) we have two charm and two anticharm jets. Due to the resonance $h^{\prime\prime}$ we expect that in experiment one should find an enhancement in the $c\bar{c}$ mass distribution at $m_{h^{\prime\prime}}$. Such a study would profit enormously if charm tagging can be realized, a challenging task at the LHC. However, it has been shown to be feasible. ATLAS, for instance, reports a charm-tagging efficiency of about 30% Aad and others (2023).
@@ 2 Results: generalities and integrated cross sections
We want to show our results for the calculation of charm-quark production, $c\bar{c}c\bar{c}$, in proton–proton collisions at the LHC. We compare the predictions of the MCPM’ versus those of the SM. We employ the MadGraph5 Monte-Carlo event generator to simulate the events Alwall et al. (2011). The MCPM has been implemented in MadGraph5 Brehmer et al. (2013). We extend this implementation by adding the effective gluon-gluon-$h^{\prime\prime}$ interaction. The $h^{\prime\prime}$ boson coupling to the gluons (and photons) is only possible via loops. The calculation of this effective interaction is presented in Appendix A.
For the process of associated $h^{\prime\prime}$ charm production we adopt for the center-of-mass energy the current LHC value of $\sqrt{s}=$ 13.6 TeV. The mass of the pseudoscalar $h^{\prime\prime}$ boson is chosen to match the possible resonance observed at CMS, that is, $m_{h^{\prime\prime}}=95.4$ GeV, (5). The total width of the $h^{\prime\prime}$ boson has been calculated in Maniatis and Nachtmann (2009) giving, for the chosen mass value
@@
For the parton densities we use the set MSHT20lo_\!\_\!as130 Bailey et al. (2021), the factorization and the renormalization scales are set to $\mu_{F}=\mu_{R}=95.4$ GeV.
@@ 3 Distributions
Now we come to the distributions for the reactions (6) and (7). The $c$ and $\bar{c}$ quarks in these reactions will manifest themselves as $c$ and $\bar{c}$ jets. Therefore, we will now specify the cuts which we apply at parton level in order to obtain distributions which should be also valid at the jet level. In the following we assume that the $c$ and $\bar{c}$ jets are identified by flavor-tagging. We are aware of the fact that this is a tough requirement for experimental analyses. But, as already mentioned, prospects to achieve this are quite good Aad and others (2023); CMS Collaboration (2023).
@@ 3.2 Cross sections for $p+p\to h^{\prime\prime}+c+\bar{c}+X$ with $h^{\prime\prime}\to c+\bar{c}$
The $h^{\prime\prime}$ boson couples in particular to the charm quarks with a coupling strength proportional to the top-quark mass. This is a special feature of the MCPM’ - eventually originating from the underlying CP symmetry; see Maniatis and Nachtmann (2025). With the application of the cuts as outlined in Section 3.1 we find a total cross section of
@@ 3.3 Cross sections for $p+p\to c+\bar{c}+c+\bar{c}+X$
Figure 9 presents a comparison of SM calculations for $d\sigma/dM_{c\bar{c}}$ obtained using different theoretical approaches. The production of the $c\bar{c}c\bar{c}$ final state within the Standard Model has been thoroughly analyzed in Maciuła and Szczurek (2013); van Hameren et al. (2014, 2015), where both single-parton scattering (SPS) and double-parton scattering (DPS) mechanisms were considered. In the present study, we adopt the same theoretical frameworks and computational setup. The corresponding numerical results here are obtained using the KaTie Monte Carlo event generator van Hameren (2018), employing both the collinear and $k_{T}$-factorization approaches [^1] We have checked numerically, that the results from MadGraph are consistent with those obtained using the KaTie generator in the collinear mode.. For the unintegrated (transverse momentum dependent) parton distribution functions (uPDFs), we use the PB-LO-2020-set1 model Jung et al. (2025), obtained within the parton-branching method Hautmann et al. (2018), as well as the widely adopted Kimber-Martin-Ryskin (KMR) model Kimber et al. (2001); Watt et al. (2003), constructed with the strong-ordering condition and by using the MSHT20lo_\!\_\!as130 collinear PDFs as input. These two models can be reliably used in the probed kinematical domain, since they are based on the DGLAP evolution equations. The use of distributions specifically designed for small values of $x$ (as is the case in most of the literature) may not be fully justified in this region. In the case of DPS, we follow a commonly used uncorrelated factorized ansatz Gaunt and Stirling (2010); Diehl et al. (2012); Blok:2011bu; Blok:2013bpa, taking as effective cross section $\sigma_{\mathrm{eff}}=15$ mb. This is a value extracted from experimental analyses of various processes at different kinematic configurations (for recent results see e.g. Aaboud and others (2016); Aaij and others (2017); Chatrchyan and others (2014); Sirunyan and others (2018b)). The $k_{T}$-factorization approach with the PB-LO-2020-set1 uPDFs yields smaller cross sections in the signal region compared to the collinear calculations (see Fig. 9(a)). In contrast, the calculations performed within the KMR uPDFs yield higher cross sections and result in distributions that are very close to those obtained in the collinear approach (see Fig. 9(b)). Moreover, as evident from the figure, the contribution from the double parton scattering (DPS) mechanism is significantly lower than that of the standard single parton scattering (SPS) process. In the signal region the DPS effect is around 15% of the SPS cross section.
@@ 4 Conclusions
In the present paper we have worked out further consequences of the model MCPM’ for experiments at the LHC. The MCPM’ is a two-Higgs-doublet model (2HDM) which has been introduced in Maniatis and Nachtmann (2025). The Higgs sector of the MCPM’contains, as in any 2HDM, three neutral Higgs bosons ($\rho^{\prime},h^{\prime},h^{\prime\prime}$) and a charged pair ($H^{\pm}$). The CP-symmetry relations forming the basis of the MCPM’ fix the couplings of these Higgs bosons to the fermions to a large extent. The $\rho^{\prime}$ couples to the fermions like the SM Higgs particle. The bosons $h^{\prime}$ and $h^{\prime\prime}$ couple mainly to charm quarks in a scalar and pseudoscalar way, respectively. The corresponding coupling constants are, however, proportional to the top-quark mass.
In Maniatis and Nachtmann (2025) the MCPM’ was compared to the results for $\gamma\gamma$ production from 30 where a possible enhancement at $M_{\gamma\gamma}=95.4$ GeV was reported. Choosing the mass of the pseudoscalar boson $h^{\prime\prime}$ of the MCPM’ at this value compatibility of the MCPM’ with the results of 30 was found.
@@
On the experimental side the identification of $c$ and $\bar{c}$ jets is improving (see Aaij and others (2015); Vértesi (2021); Sirunyan and others (2018a)). For example, the LHCb collaboration has measured the cross section for $c\bar{c}$ dijets Aaij and others (2021) (see also Maciuła et al. (2022)). The CMS and ATLAS would need to perform their own analyses including jet missidentification probabilities. In our opinion the ALICE-3 project 20 would be ideal in the future to perform experimental studies corresponding to our theoretical studies discussed here.
@@ Acknowledgements.
We now determine the effective $h^{\prime\prime}$–gluon–gluon interaction. This interaction has no tree-level contribution but arises from a triangle quark loop. Based on this interaction, we get in particular contributions to the process $p+p\to h^{\prime\prime}+c+\bar{c}+X$ via gluon–gluon fusion; see Fig. 2. We estimate that for a mass $m_{h^{\prime\prime}}\sim 95.4\penalty 10000\ \mathrm{GeV}$ in proton–proton collisions, with a center-of-mass energy of $13.6\penalty 10000\ \mathrm{TeV}$, the corresponding Bjorken-$x$ for the gluons is about $0.007$. Gluon densities in the protons are known to be very high at low Bjorken-$x$. On the other hand, the coupling of a gluon pair to the $h^{\prime\prime}$ boson is loop suppressed, with the dominant contribution arising from the charm-quark loop. In the following, we use the $\gamma$-matrix conventions of Bjorken and Drell (1965) and the Levi-Civita symbol with the normalization $\varepsilon_{0123}=1$.
@@
At one loop level we have in the MCPM’ for (A.1) the diagrams shown in Fig. 19. The evaluation of the diagrams of Fig. 19 is straightforward using the $h^{\prime\prime}q\bar{q}$ couplings from (A.81) of Appendix A of Maniatis and Nachtmann (2025):
@@
With the mass values for the quarks listed in Navas and others (2024) and $\alpha_{s}=0.12$, we find numerically:
@@
Appendix B Digression on total cross sections and comparison of $h^{\prime\prime}$ and $Z$ distributions
Now we want to calculate the total cross section for $h^{\prime\prime}$ production, composed of the cross sections for the Drell–Yan process with $c\bar{c}$ fusion and the $GG$ fusion process discussed in (37) and (38) of Maniatis and Nachtmann (2025) and the process (6) here. For this we shall first make some general considerations.